a man is 3 yrs times older than his son. 4 yrs ago he was 4 times as his son was at that time. how old are they now?
Answer by ptfile(81)

A man is 3 yrs times older than his son.
m=3s
4 yrs ago he was 4 times as his son was at that time.
3s-4=4(s-4)
3s-4=4s-16
12=s
36=m
His current age is 36 years old while his son current age is 12 years old.